OpenStudy (anonymous):

Triangle DEF where angle E is a right angle. DE measures x; DF measures 55; angle F measures 49 degrees. In △DEF, solve for x. Round the solution to the nearest hundredth.

OpenStudy (anonymous):

Solve for x knowing that \[ x^2 + 55^2 = EF^2\\ EF \cos(49) = 55\\ \]